OpenStudy (anonymous):

Can someone please answer this question, 4^3x+1 = 7^x thanks.

OpenStudy (anonymous):

Are you asking for solving the solution \[ 4^{3x}+1=7^x \] ? Because your notation is somewhat ambiguous.

OpenStudy (anonymous):

*solving the equation

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

actually the +1 is part of the exponent

OpenStudy (anonymous):

Okay because the other one has no real solutions ;).

OpenStudy (anonymous):

The solution is \[ -2\,{\frac {\ln \left( 2 \right) }{6\,\ln \left( 2 \right) -\ln \left( 7 \right) }} \]

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

Oh ic, any chance you could show me your steps?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yea thats how to do it, but you messed up somewhere I guess.

OpenStudy (anonymous):

The main idea is take the log on both sides and then use logarithm laws.

OpenStudy (anonymous):

Alright thanks to you both for your help, I appreciate it and the answer says -ln4 / 3ln4 - ln7

jimthompson5910 (jim_thompson5910):

\[\Large 4^{3x+1} = 7^{x}\] \[\Large \ln(4^{3x+1}) = \ln(7^{x})\] \[\Large (3x+1)\ln(4) = x\ln(7)\] \[\Large 3x\ln(4)+\ln(4) = x\ln(7)\] \[\Large \ln(4) = x\ln(7)-3x\ln(4)\] \[\Large \ln(4) = x\left(\ln(7)-3\ln(4)\right)\] \[\Large \frac{\ln(4)}{\ln(7)-3\ln(4)}=x\] \[\Large x=\frac{\ln(4)}{\ln(7)-3\ln(4)}\]

myininaya (myininaya):

something happen to your 3 jim

jimthompson5910 (jim_thompson5910):

yep it ran away

myininaya (myininaya):

ok you got it lol

OpenStudy (anonymous):

Yea my answer was the same, but it looks different because I let a CAS do it :D

OpenStudy (anonymous):

oh lol, thanks for the help!

jimthompson5910 (jim_thompson5910):

to match what the book has, follow the steps given below \[\Large x=\frac{\ln(4)}{\ln(7)-3\ln(4)}\] \[\Large x=\frac{\ln(4)}{-3\ln(4)+\ln(7)}\] \[\Large x=\frac{\ln(4)}{-(3\ln(4)-\ln(7))}\] \[\Large x=\frac{-\ln(4)}{3\ln(4)-ln(7)}\]

jimthompson5910 (jim_thompson5910):

either way, it's the same answer

OpenStudy (anonymous):

Oh alright, makes sense , thanks

OpenStudy (anonymous):

those threes are tricky